OpenStudy (anonymous):

what is the probability p(0

OpenStudy (anonymous):

@grabill @jigglypuff314

OpenStudy (anonymous):

@DarrenMadx

OpenStudy (michele_laino):

the sum of probabilities is: \(0.3+0.5+0.15+0.15+0.25=1.35\) so the requested probability, regards the probabilities for events \(1\) and \(2\) only: So we can write: \[p = \frac{{0.3 + 0.5}}{{1.35}} = ...?\]